OpenStudy (anonymous):

Simplify this: 4√3 ---------- √5 - 3√2 = 4√3 x √5 + 3√2 ------ ---------- √5 - 3 √2 √5 + 3 √2 Basically I am not sure why there is a positive in the second fraction. Shouldn't the denominator be √5 - 3 √2 for both fractions? Thanks.

OpenStudy (mertsj):

\[\frac{4\sqrt{3}}{\sqrt{5}-3\sqrt{2}}\times\frac{\sqrt{5}+3\sqrt{2}}{\sqrt{5}+3\sqrt{2}}=\frac{4\sqrt{15}+12\sqrt{6}}{5-18}=\frac{4\sqrt{15}+12\sqrt{6}}{-13}\]

OpenStudy (mertsj):

The denominator is correct because you multiply by the conjugate to eliminate the radical from the denominator.

OpenStudy (anonymous):

In order to get rid of sqrt under denominator, apply (a -b) (a +b) = a^2 - b^2: 4√3 /( √5 + 3√2) = [ 4√3 ( √5 + 3√2) ] /[ 5 - 18 ] = - ( 4√15 + 12√6 ) / 13

OpenStudy (anonymous):

Just attempt to make it more presentable: = [ 4√3 ( √5 + 3√2) ] /[ 5 - 18 ] = - ( 4√15 + 12√6 ) / 13